OpenStudy (anonymous):

help with this question: solve with elimination. 6x+y=-10 3x-y=3

OpenStudy (anonymous):

add both equations

OpenStudy (anonymous):

add them 6x+y=-10 +)3x-y=3 ------------- 9x=-7 x=-7/9